Please help me solve this problem: 0.5x - 0.2y = 1
Answer by Fombitz(32388)

Are you trying to solve for y as a function of x?
0.5x - 0.2y = 1
Multiply by 10.
5x-2y=10
-2y=-5x+10
y=(5/2)x-5
That's all you can do.